# Multi-Cell Decoder and Mutual Learning for Table Structure and Character Recognition

### v1

@@ 2 Related Work
Table structure recognition has been studied for a long time, and approaches based on hand-crafted features and heuristic rules [10, 12, 30] were proposed, but their application was limited to simple tables or tables with predefined patterns. With the development of deep learning, methods that automatically learn table structural patterns [36, 25, 24, 26, 28] have become mainstream. These studies can be divided into approaches based on object detection and segmentation [28, 25], and approaches based on sequential token prediction models [20, 33].
@@
For the sequential token prediction approaches, a simple image caption model can be utilized for cell detection because the order of cells is uniquely determined. Ye et al. [33] and Nassar et al. [20] proposed Transformer models with two types of decoders for table structure recognition and cell localization.
In the 2020s, researchers are investigating end-to-end models that learn both table structure and cell content recognition tasks [3]. Zhong et al. [39] proposed a model that uses a ResNet [6] encoder and two LSTM [8] decoders to recognize both table structure and cell contents, but its performance is inferior to solutions using external OCR.
@@ 3.2 Decoder
In practice, the attention layer is divided into several groups, each of which pays attention independently and combines the outputs at the end. This is called multi-headed attention [29]. Through the above mechanism, the Transformer can focus on specific values of $Y$ and incorporate them into the $X$ series.
@@ 5.1.3 PubTabNet250
Ly and Takasu [17] extracted tables with 250 or more HTML tokens from PubTabNet and created a subset named PubTabNet250. This subset was utilized originally [17] to demonstrate the effectiveness of the local attention mechanism. We also used this subset to conduct ablation studies in Section 5.4.
@@ 5.3.1 FinTabNet
We evaluated the experimental results of structure recognition and total recognition using the TEDS metric. Table 5.3.1 compares the TEDS scores in the test set between the proposal and previous models. The proposal outperforms all previous models with scores of 98.87% and 97.69%.
@@ 5.3.2 PubTabNet
We evaluated the experimental results of table recognition on the validation set using the TEDS metric. Table 5.3.2 compares the scores between the proposal and previous methods. The proposal outperforms all previous methods with scores of 98.16%, 95.53% and 96.87% on simple tables, complex tables, and all tables, respectively. We also evaluated the proposal on the ICDAR evaluation set. Table 5.3.2 compares the TEDS score of the proposed method with the top 10 solutions of the ICDAR competition [34].
@@ 5.4 Ablation Studies
We conducted additional experiments for ablation studies using PubTabNet250 dataset for training and evaluation.
@@ 5.4.1 Effectiveness of Multi-Cell Decoder and Mutual Learning
Since the previous study [17] focused on performance for long tables, we also calculated scores for tables with more than 500, 600, and 700 structure tokens. The MC decoder significantly outperforms the baselines at all table lengths, and BML further improves table recognition performance.
@@ 5.4.2 Window Size of Cell Decoder
The effect of window size is not as clear as the evaluation results [17] for the HTML decoder, and the change may be explained by randomness. Consequently, we decided to use the window size of 300 as the optimal value.

### v2

# https://researchmap.jp/t.kat

@@ 2 Related Work
Table structure recognition has been studied for a long time, and approaches based on hand-crafted features and heuristic rules [11, 13, 32] were proposed, but their application was limited to simple tables or tables with predefined patterns. With the development of deep learning, methods that automatically learn table structural patterns [39, 27, 26, 28, 30] have become mainstream. These studies can be divided into approaches based on object detection and segmentation [30, 27], and approaches based on sequential token prediction [21, 36].
@@
For the sequential token prediction approaches, a simple image caption model can be utilized for cell detection because the order of cells is uniquely determined. Ye et al. [36] and Nassar et al. [21] proposed Transformer models with two types of decoders for table structure recognition and cell localization. Peng et al. [25] achieved performance comparable to a model using a deep convolutional encoder while significantly reducing parameters by introducing a convolutional stem.
In the 2020s, researchers are investigating end-to-end models that learn both table structure and cell content recognition tasks [3]. Zhong et al. [42] proposed a model that uses a ResNet [6] encoder and two LSTM [8] decoders to recognize both table structure and cell contents, but its performance was inferior to models using external OCR.
@@
TabRecSet [35] is a bilingual dataset containing rotated and distorted tables in real photographs for three tasks, namely table detection, structure recognition, and cell content recognition. Detection of such tables is outside the scope of this paper.
@@ 3.2 Decoder
In practice, the attention layer is divided into several groups, each of which pays attention independently and combines the outputs at the end. This is called multi-head attention [31]. Through the above mechanism, Transformer can focus on specific values of $Y$ and incorporate them into the $X$ series.
@@ 5.1.3 PubTabNet250
Ly and Takasu [18] extracted tables with 250 or more HTML tokens from PubTabNet and created a subset named PubTabNet250. They also introduced subsets for tables containing at least 500, 600, and 700 tokens. These subsets were utilized originally [18] to demonstrate the effectiveness of the local attention mechanism. We also utilized these subsets to conduct ablation studies in Section 5.4, approximately reducing training time from 179 hours to 45 hours per model.
@@ 5.3.1 FinTabNet
We evaluated the experimental results of structure recognition and total recognition using the TEDS metric. Table 5.3.1 compares the TEDS scores in the test set between the proposal and previous models. The proposal outperforms the previous models with scores of 98.87% and 97.69%. The inference time using the 4 GPUs was 3.78 hours.
The total TEDS score of the proposal was lower than the score of VAST [10], which could be explained by the fact that VAST exploits external OCR for cell content recognition. In contrast, the structural TEDS score of VAST was lower than those of end-to-end approaches [20, 19, 18], including the proposal.
@@ 5.3.2 PubTabNet
We evaluated the experimental results of table recognition on the validation set using the TEDS metric. Table 5.3.2 compares the scores between the proposal and previous methods. The proposal outperforms all previous methods with scores of 98.16%, 95.53% and 96.87% on simple tables, complex tables, and all tables, respectively. The inference time using the 4 GPUs was 3.23 hours.
We also evaluated our proposal on the evaluation set. Table 5.3.2 compares the scores of the proposal with the top 10 solutions of the ICDAR competition [37]. The high scores achieved on both sets indicate high generalization performance of the proposal. The inference time was 3.13 hours.
The score of the proposal was higher than the score of VAST [10]. PubTabNet contains a large amount of training data, and the proposed model appears to be well trained for cell content recognition tasks.
@@ 5.4 Ablation Studies
We conducted additional experiments for ablation studies using PubTabNet250 dataset for training and PubTabNet subsets for evaluation.
@@ 5.4.1 Effectiveness of Multi-Cell Decoder and Mutual Learning
Since the previous study [18] focused on performance for long tables, we also calculated TEDS scores for tables containing at least 500, 600, and 700 structure tokens. The MC decoder outperforms the baselines at all table lengths, and BML further improves table recognition performance.
The effect of BML was unclear in the structural TEDS scores but evident in the total TEDS scores. BML may still have improved the performance of implicit structure recognition and may have made an impact on cell content recognition, which requires precise content locations.
@@ 5.4.2 Window Size of Cell Decoder
In general, a window size of 300 achieved the highest score, with the exception of tables containing more than 500 tokens, where a window size of 100 achieved the highest score. Tables with many cells tend to have fewer characters per cell, and a shorter window may be sufficient.
It should be noted that we used the PubTabNet250 dataset for training, and the performance for tables with fewer structure tokens was lower than the scores in Table 5.3.2. We selected the window size of 300 as the optimal value for the entire PubTabNet dataset containing tables with fewer tokens from the perspective of generalization performance.